OpenStudy (anonymous):

If k is a real number, what is the value of k if the following equation only has 1 solution? kx^2+36x+k=0

OpenStudy (anonymous):

It will help to divide the equation by "k" and complete the square.

OpenStudy (anonymous):

how would you complete the square?

OpenStudy (anonymous):

no need

OpenStudy (anonymous):

just check wether b^2-4ac =0

OpenStudy (anonymous):

a=k, b=36, and c=k

OpenStudy (anonymous):

(36)^2-4k^2=0 k^2=9^2 k=9

OpenStudy (anonymous):

sorry k=18

OpenStudy (anonymous):

or k=-18

OpenStudy (anonymous):

sorry, i just dont follow-what do you do after 36^2-4k^2?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

36^2-4k^2=0 4k^2=36^2 k^2=(36^2)/(4) take the square root on both sides |k|=36/2=18 Then, k=+18 or k=-18

OpenStudy (anonymous):

Oh! I get it now. thank you so much!